ENCAPSULATION OF WATER MOLECULES BY DIBENZO-18-CROWN-6-ETHER IN A SUPERSONIC JET

\underline{R. KUSAKA}, Y. INOKUCHI, K. KAWASAKI AND T. EBATA, Department of Chemistry, Graduate School of Science, Hiroshima University, Higashi-Hiroshima, 739-8526, Japan.

UV and IR spectra of jet-cooled Dibenzo-18-crown-6-ether (DB18C6) and DB18C6-(H_{2}O)_{n} are observed by laser induced fluorescence (LIF), UV-UV hole-burning and IR-UV double resonance spectroscopy ^{a}. The LIF spectrum exhibited complicated vibronic structure in the region of 35500-36800 cm^{−1}. The UV-UV hole-burning and IR-UV double resonance spectra suggest the existence of two conformers for DB18C6; one is ”boat-type” structure having C_{2}v symmetry, and the other is ”chair-type” structure having C_{i} symmetry. For the hydrated clusters, we assigned most of them to the ”boat-type” DB18C6-(H_{2}O)_{n=1−4}. Details of the encapsulation structure in DB18C6-(H_{2}O)_{n=1−4} will be discussed by the comparison of the observed IR spectra with those obtained by quantum chemical calculation.

aR. Kusaka, Y. Inokuchi, and T. Ebata, Phys. Chem. Chem. Phys. 9, 4452 (2007).